# Escalating LLM-based Code Translation Benchmarking into the Class-level Era

### v2

@@ 6. Implications
Implications for researchers: As the first study focusing on class-level code translation, this paper revealed the defects of previous method-level code translation benchmarks and exposed LLMs’ inability to translate code for an entire class in RQ1. In particular, our contributed benchmark with class-level code, diverse dependencies, and alignment with practical development significantly fosters the code translation research to the era of class level. In addition, RQ4 unveiled a series of failure cases that the best-performing LLM made, enlightening future research constantly focusing on resolving syntax, functional, and runtime errors that LLMs also suffer in method-level code translations, but also pay attention to those errors mainly occur in class-level code translations, such as library-related, function/variable usage, and code consistency/completeness issues.
Implications for practitioners: Based on our contributed benchmark, namely ClassEval-T, we conducted extensive experiments and concluded insightful findings, which can guide developers’ practical usage of LLMs for class-level code translation. For example, RQ2-3 uncovered the discrepancies in correctness and dependency awareness among diverse translation strategies, LLMs, and PLs, suggesting that powerful LLMs, such as DeepSeek, and Python-oriented translations are more suitable for holistic strategy; The selection of translation strategy for smaller LLMs depends on different scenarios, and C++-oriented translations prefer the min-dependency strategy. Therefore, choosing a suitable translation strategy can boost the class-level code translation performance. Besides, RQ4 also reported and categorized diverse errors the best-performing LLMs may make, precisely reminding practitioners to notice their potential risks when deployment.
@@ 7. Threats to Validity
Threats to external validity relates to the generality of our experimental findings and failure taxonomy (Zhang et al., 2024; Liu et al., 2024c). Our newly constructed benchmark specifically targets code translation among Python, Java, and C++ without the consideration of many other domain-specific PLs. Nonetheless, ClassEval-T contains more practical coding tasks on the class level than any other previous counterparts, making our findings and failures of LLMs more useful and valuable for daily development. Besides, Python, Java, and C++ are the most widely studied and used PLs, which are much more meaningful for exploration. In the future, we will extend ClassEval-T to other domain-specific PLs and investigate their code translation performance. As for failure taxonomy, we thoroughly evaluated 1,397 class-level error cases made by DeepSeek across all three translation strategies. Besides, a thematic-analysis-driven methodology is strictly followed, ensuring the generality and correctness of our conclusion. Therefore, we believe the above threats are minimal.
Threats to internal validity lies in the potential of data leakage and processing involving manual handling (e.g., benchmark manual construction and manual evaluation) (Ouyang et al., 2024; Xia et al., 2023). ClassEval-T is a manually constructed code translation benchmark extended from ClassEval with extra parallel Java and C++ programs. It is inevitable that Python programs in ClassEval may have been crawled by our studied LLMs for training, but their corresponding Java and C++ programs are unseen to LLMs. No matter which translation pairs are experimented with, at least code samples on one side (source or target) are assured unseen to LLMs. In addition, 0% of translated samples exactly match the ground truths. Therefore, the data leakage threat is extremely limited. To guarantee the quality of our constructed benchmark, we predefine rigorous translation principles and include multiple participators to alleviate the subjectiveness and mistakes in manual coding. As for the accuracy of manual evaluation in dependency analysis (RQ3) and failure analysis (RQ4), multiple experienced participators are also included for assessment. Moreover, we follow the methodology of thematic analysis with a detailed codebook for guidelines and a rigorous, double-blind, and dual-review process in failure categorization, thereby minimizing this internal threat.

### v3

@@ 6. Implications
Implications for researchers: As the first study focusing on class-level code translation, this paper revealed the defects of previous method-level code translation benchmarks and exposed LLMs’ inability to translate code for an entire class in RQ1. In particular, our contributed benchmark with class-level code, diverse dependencies, and alignment with practical development significantly fosters the code translation research to the era of class level. While repository-level code translation remains a long-term objective, significant challenges arise from cross-file dependencies and the context window limitations of LLMs, making complete translation in a single pass difficult. However, experimental findings indicate that holistic translation is optimal when the context window is sufficiently large. Therefore, future repo-level research should aim to maximize granularity in code splitting to leverage larger context windows effectively. This study suggests that, for subsequent repo-level investigations, adopting larger code units such as classes or modules can significantly improve translation quality, providing a crucial stepping stone towards achieving comprehensive repository-level translations. Our findings indicate that, even at class-level granularity, LLMs struggle with structural integrity, dependency handling, and code correctness, highlighting the necessity of a phased approach. By first addressing the challenges at the class level, our work provides a solid foundation, offering a finer-grained perspective that positions future research to progressively and efficiently tackle real-world translation scenarios. In addition, RQ4 unveiled a series of failure cases that the best-performing LLM made, enlightening future research constantly focusing on resolving syntax, functional, and runtime errors that LLMs also suffer in method-level code translations, but also pay attention to those errors mainly occur in class-level code translations, such as library-related, function/variable usage, and code consistency/completeness issues.
Implications for practitioners: Based on our contributed benchmark, namely ClassEval-T, we conducted extensive experiments and concluded insightful findings, which can guide developers’ practical usage of LLMs for class-level code translation. For example, RQ2-3 uncovered the discrepancies in correctness and dependency awareness among diverse translation strategies, LLMs, and PLs, suggesting that powerful LLMs, such as DeepSeek, and Python/Java-oriented translations are more suitable for holistic strategy; The selection of translation strategy for smaller LLMs depends on different scenarios, and for C++-oriented translations, the min-dependency strategy and holistic approach are neck-to-neck in terms of performance. Therefore, choosing a suitable translation strategy can boost the class-level code translation performance. Besides, RQ4 also reported and categorized diverse errors the best-performing LLMs may make, precisely reminding practitioners to notice their potential risks when deployment.
@@ 7. Threats to Validity
Threats to external validity relates to the generality of our experimental findings and failure taxonomy (Zhang et al., 2024; Liu et al., 2024c). Our newly constructed benchmark specifically targets code translation among Python, Java, and C++ without the consideration of many other domain-specific PLs. Nonetheless, ClassEval-T contains more practical coding tasks on the class level than any other previous counterparts, making our findings and failures of LLMs more useful and valuable for daily development. Besides, Python, Java, and C++ are the most widely studied and used PLs, which are much more meaningful for exploration. As for failure taxonomy, we thoroughly evaluated 1,243 class-level error cases made by DeepSeek across all three translation strategies. Besides, a thematic-analysis-driven methodology is strictly followed, ensuring the generality and correctness of our conclusion. Therefore, we believe the above threats are minimal.
Threats to internal validity lies in the potential of data leakage and processing involving manual handling (e.g., benchmark manual construction and manual evaluation) (Ouyang et al., 2024; Xia et al., 2023). ClassEval-T is a manually constructed code translation benchmark extended from ClassEval with extra parallel Java and C++ programs. It is inevitable that Python programs in ClassEval may have been crawled by our studied LLMs for training, but their corresponding Java and C++ programs are unseen to LLMs. No matter which translation pairs are experimented with, at least code samples on one side (source or target) are assured unseen to LLMs. In addition, 0% of translated samples exactly match the ground truths. Therefore, the data leakage threat is extremely limited. A related concern is the potential future contamination of open-source benchmarks as LLMs evolve and may eventually be trained on ClassEval-T. To mitigate this, we intend to periodically update the dataset using program mutation rules that systematically transform benchmark programs while preserving functional equivalence. This approach maintains benchmark reliability, ensuring its continued relevance and reducing the risk of LLMs memorizing specific samples. Additionally, to guarantee the quality of our constructed benchmark, we predefine rigorous translation principles and include multiple participators to alleviate the subjectiveness and mistakes in manual coding. As for the accuracy of manual evaluation in dependency analysis (RQ3) and failure analysis (RQ4), multiple experienced participators are also included for assessment. Moreover, we follow the methodology of thematic analysis with a detailed codebook for guidelines and a rigorous, double-blind, and dual-review process in failure categorization, thereby minimizing this internal threat.
@@ 8. Conclusion and Future Work
Based on our research findings, our future work aims to enhance the performance of LLMs in class-level code translation tasks by integrating structural information (e.g., AST), to minimize syntax errors and improve translation accuracy. Additionally, we will extend ClassEval-T to include a broader range of PL pairs, particularly those with strong industrial relevance, such as C++ to Rust.